OpenStudy (anonymous):

Which expression is equivalent to 4^7 * 4^-5 ? A) 4^12 B) 4^2 C) 4^-2 D) 4^-35 I got B as my answer am I correct ?

OpenStudy (alexandervonhumboldt2):

when we multiply, we add the exponents: \[\huge\bf a^b*a^c=a^{b+c}\]

OpenStudy (alexandervonhumboldt2):

so we have: \[\huge\rm 4^7*4^{-5}=4^{7+(-5)}=?\]

OpenStudy (anonymous):

2

OpenStudy (alexandervonhumboldt2):

yeah: so we have: \[4^2\]

OpenStudy (anonymous):

Alrighty Thank you :)